Use only the ultimate-strength design (strength design and calculate the design strength (ultimate strength capacity Mn) o
Answer #1

Following steps are involed in the solution as shown in images

to 2017 he 29 28 4#11 3 K 12 Soin A# = 6.25 in² t = 4700 psi = 4.7 ksi fy - 75000 þsi b=20 bw = 12 d - 29 = 75ksi 1. SAflange = 20x4 = 80 in Ac 7 A flange not toue. . a hf Hence assumption is Hence strain/ stress profile will change. 70.003C₂ = Tz >> 0.85 fcabw = (As-Asp).fy. 0.85% 4.7xax)2 = ( 6-25 -1.705)x75 a = 7oll in -1 jd₂= de a = 29 - 7.11 - 25.445 2 j